One canned juice drink is 15% orange juice;another is 10% orange juice.How many liters of each should be mixed together in order to get 5L that is 12% orange juice?
Answer by stanbon(75887)

One canned juice drink is 15% orange juice;another is 10% orange juice.How many liters of each should be mixed together in order to get 5L that is 12% orange juice?
-----
Equation:
orange + orange = orange
0.15x + 0.10(5-x) = 0.12*5
-----
Multiply thru by 100 to get:
15x + 10*5 - 10x = 12*5
----
5x = 2*5
x = 2L (amt. of 15% juice needed)
5-x = 3L (amt. of 10% juice needed)
------
Cheers,
Stan H.
==============